We Do Bankruptcies is a law firm that specializes in providing advice to firms in financial distress. It prospers in recessions when other firms are struggling. Consequently, its beta is negative, −0.2.

a. If the interest rate on Treasury bills is 5% and the expected return on the market portfolio is 20%, what is the expected return on the shares of the law firm according to the CAPM? (Enter your answer as a whole percent.)

 

b. Suppose you invested 90% of your wealth in the market portfolio and the remainder of your wealth in the shares in the law firm. What would be the beta of your portfolio? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a. Expected return = rf + b(rm - rf) = 0.05 + (-0.2)(0.20 - 0.05) = 0.02 or 2%

b. Portfolio beta = (0.70 × Bmarket) + (0.30 × Blaw firm)

Portfolio beta = (0.90 × 1.0) + [0.10 × (-0.2)]

Portfolio beta = 0.88